OpenStudy (anonymous):

A pump can throw a stream of water up to 29.6 m. What is the initial speed of the stream when it leaves a hoze nozzle?

OpenStudy (anonymous):

Loss of kinetic energy = gain in potential energy (1/2) m v^2 = m g h v = sqrt(2 g h) take it from there.

OpenStudy (cail):

KE = PE KE = (V^2)/2 PE = GH (Ignoring mass) Solve for H given V =29.6 and G = 9.8

OpenStudy (mrnood):

Cail - the question gives H= 29.6m from which the KE can be calculated (the equations are above - you can 'ignore mass' since it occurs on both sides of the equation). As Douglas says - you then solve for V given that initial KE = final PE